OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y - 4 = 3(x - 2) y - 2 = 3(x - 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (anonymous):

Equation of the given line: \[\large y - y_1 = m(x-x_1)\] So, y - 4 = 3(x-2)